the length of rectangle is four times its width. the area of the rectangle is 676 square feet, find the dimensions of the rectangle.

Answer by checkley79(3341)

L=4W
AREA=LW
676=4W*W
676=4W^2
W^2=676/4
W^2=169
W=SQRT169
W=13 ANS FOR THE WIDTH.
L=4*13=52 ANS FOR THE LENGTH.
PROOF:
676=13*52
676=676